OpenStudy (anonymous):

The sum of 4 times a number and is -13

OpenStudy (anonymous):

?

OpenStudy (anonymous):

3/12 ×2=11

OpenStudy (anonymous):

where is the question?

OpenStudy (anonymous):

What is the sum of 4 times a number and 3 is -13

OpenStudy (anonymous):

im lost...

OpenStudy (anonymous):

What about 3/12× + 2 = 11

OpenStudy (anonymous):

better

OpenStudy (anonymous):

subtract 2 from both sides.

OpenStudy (anonymous):

then divide both sides by (3/12) or .25

OpenStudy (anonymous):



OpenStudy (anonymous):

3.5 + z= 3.7